

********************************************************************
UARS WEEKLY STATUS REPORT - 15 OCT 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods. Due to a mis-communication with
the MPG, there were no stellar calibration experiments executed for
six days this week.
***********
SUSIM
***********
SUSIM successfully performed a "six monthly" calibration involving three
of its four deuterium lamps. In addition, SUSIM gathered its normal
solar, offset and occultation data this past week without apparent
problems.
We have requested that the SUSIM level 3 irradiances through UARS day
2523 be sent to the GSFC DAAC.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers. The AXIS activated in its half power mode on October 3.
Since then, AXIS has been producing energy deposition from 8 of 16 channels
from precipitating electrons derived from measured X-rays.
The FOT and the PEM team are updating the PEM activation sequence which
effects the PEM/HEPS (> 30 keV particles). Current activation sequence
reduces the activation time by about 1 minute. Additional analization shows
that the activation sequence may be reduced by about another minute.
Current activation takes about 7.5 min for full activation. During the
activation sequence, various components of PEM become active at all different
times, beginning with the PEM magnetometer (at about on activation time) and
ending with the HEPS-2 components (7.5 min later).
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating range. HALOE science data products continue to be
excellent.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
The HRDI instrument status telemetry showed some anomalous behavior this week,
beginning on Oct 7. At the time of this writing, the cause of this anomaly
and the current status of the telemetry are still under investigation. Only
housekeeping TM was affected; the HRDI instrument continues to collect
and report valid science data.
The HRDI team wishes the above news was the worst to be reported this
week, but it is not. It is with great sadness that we must notify the
rest of the UARS family of the most untimely death of one of our own.
Dr. Mark Burrage passed away suddenly this Sunday, October 10, at the
age of 39. A graduate of the University of Leicester, England, Mark
came to the Space Physics Research Lab in the Department of
Atmospheric, Oceanic, and Space Sciences (AOSS) at the University of
Michigan as a postdoc in 1988 where he first worked on analysis of data
from the Atmospheric Explorer missions. He soon became involved in the
HRDI project and in 1990 he joined the SPRL research faculty, where in
1996 he was promoted to the level of Associate Research Scientist. His
work was recognized by the AOSS Department's Excellence in Research
Award in 1996, and in 1998 he was honored with the University of
Michigan's Research Scientist Recognition Award. In 1996 Mark joined
the Michigan Aerospace Corporation, a new venture founded by Paul Hays,
where he continued to pursue research and development of new
space-based instrumentation. His career was cut short in midstream,
but not before he made significant contributions to our understanding
of the aeronomy and dynamics of the middle atmosphere, with over 40
peer-reviewed publications to his name. Mark's work is best
exemplified by his analyses of HRDI mesospheric winds and airglow to
discern the long-term climatological signal of the migrating tides.
His tenacious and meticulous scientific style led him to discover the
signature of the mesospheric quasibiennial oscillation (MQBO) in HRDI
and MF radar winds. He will also be remembered with gratitude by many
researchers for his generous assistance in their use and interpretion
of the HRDI data set. We in the HRDI group especially appreciate his
dedicted pursuit of the often thorny, but not always rewarding, work on
data validation. Mark is survived by his wife Margo and daughters
Katie, 6, and Rachel, 4. Memorial gifts to a trust fund for the
daughters' education are being accepted; anyone wishing to contribute
may contact Deborah Eddy for the particulars ( dkeddy@umich.edu ).
***********
WINDII
***********
WINDII is currently on, and operating normally. It is scheduled
to be on until October 23.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
8 days CLA (7.0) PROD (L1-L3) (473R-307C-17F=149L)
6 days GPI (4.05) PROD (CLAES L3B) (191R-170C-1F=20L)
28 days SUSIM (7.1) PROD (L1-L2) (49R-49C=0L)
1 day SEN (5.0) PROD (1992 DATA) (1R-1C=0L)
The CDHF is testing the HALOE V4.82 formal software delivery.
The CDHF will load August 1999 current day DAAC data in October 1999.
SUSIM has requested UARS DAYS 31-2523, Level 3BS (output V20) be sent
to the DAAC. Plans are to start previewing and loading data into DAAC
Transfer Tables. DAAC has been notified.
REHOST
- Attitude V7.3 testing was successful after the UCSS re-link. Differences
and file verification were done. The software was promoted to production
mode and UARS DAY 1534 was processed.
- Plans are to test HALOE V4.82 formal software delivery.
***********
FOT
***********
REPORTING PERIOD
This report covers 2 October, 1999 (Orbit 44031, GMT Day 275) through 8 October
1999 (Orbit 44135, GMT Day 281).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 276/0024 44046 AXIS On
The beta angle ranged between the angles of 25.7 degrees to 45.9 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
50 degrees on 12 October 1999 (DOY 285).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.7 to 1.0 degrees C. Additional battery performance data for each
battery on 01 October 1999 and 08 October 1999 are provided in Appendix D.
The clock error ranged from 11.3 to -11.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 03
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% PEM = 01
% SUSIM = 01
% FOT = 03
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 02
Zero (0) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/263 - 2287.496947
% Transponder B: 99/265 - 2287.498700
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/280 show the observatory data loss to be 24 hours, 55
minutes, 37.108 seconds (an increase of 0.384 seconds since the last report
period). This is a 0.0354 percent data loss which equals a 99.9646 percent
data capture for the mission.
The increase of 0.384 seconds occurred over one (1) days:
7 0.384 seconds on Day 276 due to an Data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The Leonid plan was completed by updating the previous years plan without the
use of the Earth Sensor. After consulting with FDF and NOAA analysts, the
impact by the Leonids to UARS is believed to be significantly less substantial
than last year. This opinion was presented at the UARS monthly and mutual
concurrence was established to make Leonid plans with the use of the Earth
Sensor.
YEAR 2000 OPERABILITY SUPPORT
The FOT received the remainder of the FDF-supplied planning aids necessary to
accommodate Y2K end-to-end testing. This included a set of 8-week and 3-day
files spanning the 1999-to-2000 rollover, as well as a set spanning the 2000-
to-2001 rollover. All of the FDF files were processed on the CMS, and
appropriate daily mission planning procedures were executed to create stored
command (ATC) loads for each rollover. As requested, the 2000-to-2001
processing was performed in a "time shifted" environment, with the UCMS system
clock set to 31 December, 2000. The FOT noticed no Y2K related errors in any
of the processing (for either rollover). The resulting ATC loads were uplinked
to the UTTS.
On Thursday we retested the first rollover for Y2K. The test again went as
planned and nothing out of the ordinary took place. We did not receive a clean
dump on the ATC but this is a known problem with the simulator and is not a Y2K
related issue.
MEETINGS
The FOT held a meeting on Wednesday, October 6th to discuss the Power and
Safehold status of UARS. Goddard Power and Attitude Engineers, Subcontractors
and the FOT were present. All attendees agreed that an Attitude Safehold must
be put place such that to minimize the amount of current driven to the
batteries at a particular instance. The difficulty lies in the area of an OBC
halt which is an analog/hardware mode that relies on a preset configuration
within the Safehold Electronics. This mode does not lend itself to
versatility. All agreed that with the OBC running a stable attitude safehold
could be obtained. The meeting ended with every body given the action to
research and consider any new options available. There were many positive
ideas brought out during this meeting. A second meeting will be held in one
month to continue along these lines. It is the general consensus that some
configuration should be in place prior to Y2k.
The UARS monthly meeting was held on Wednesday October 6th hosted by Paul
Ondrus. A briefing of the Attitude and Power Safehold meeting was given. In
addition it was decided that for the upcoming Leonids UARS would not configure
for Earth pointing safehold. The Earth Sensor has limited capability and is
used during special operations and safehold recovery.
FUTURE SIGNIFICANT EVENTS
A second meeting to discuss the UARS Safehold configuration and concerns for
the MPS and its batteries is planned for Friday, November 5th. This meeting
will be a closed meeting with only the FOT and specialized engineering support
attending.
PEM requested to have the timing in their turn on sequence tightened up in yet
another area. This is expected to be accomplished within the next few weeks.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 8 October 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
44046 276/0024 99-329 PEM PEM AXIS ON 14 WATTS
44064 277/0535 99-330 OBC Clk Adj fr 0.00 to 20.6 msec/day
44091 279/0025 99-331 OBC Clk Adj fr 20.6 to 0.00 msec/day
44100 279/1430 99-332 SUSIM Exception Clear
44125 281/0641 99-333 OBC Clk Adj fr 0.00 to 20.6 msec/day
___________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 8 October 1999
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI HRDI PTS Pulse Bus Current 44041, 44086-89
NBRECBI Y-HI NBTR B Current 44041
HRPSEQBIMON Y-HI HRDI PSE Quiet Bus Current 44088
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_____________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 8 October 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
44064 277/05:30 (04 Oct 99) from 00.0 to 20.6 msec/day
44091 279/00:39 (06 Oct 99) from 20.6 to 00.0 msec/day
44125 281/06:47 (08 Oct 99) from 00.0 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 8 October 1999
24 Sept 1999 - GMT Day 267 Beta = 1.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- --------- ----- ------- ------
1 -112.0/-134.4 -2.73/-3.99 +0.0/+0.0 16.5 0.000 0.0
2 +16.8/-67.2 +6.44/+5.15 +30.4/-15.2 24.5 1.023 22.3
3 +22.4/-56.0 +5.47/+3.87 +30.0/-16.0 24.3 1.024 22.1
08 Oct 1999 - GMT Day 281 Beta = 45.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- --------- ----- ------- ------
1 -112.0/-134.4 -3.05/-3.99 +0.0/+0.0 16.5 0.000 0.0
2 +22.4/-33.6 +5.79/+4.51 +27.6/-15.6 25.3 1.022 21.4
3 +22.4/-44.8 +4.83/+3.87 +27.2/-15.2 25.1 1.025 21.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 8 October 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-062 FOT 44085 278/1425 21435 TAC 8 Crashed
99-063 FOT 44088 278/2059 N/A TAC 8 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-061 FOT 43970 270/2216 21402 NEGATIVE ACQ.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
SEVNTFW99/41